OpenStudy (anonymous):

Verify the identity: cos^2x -sin^2x=1-2sin^2x Is this cos(2x)=cos(2x) ???

OpenStudy (anonymous):

cos^2x = 1-sin^2x. 1-sin^2x - sin^2x = 1 - 2sin^2x. hence cos^2x-sin^2x = 1-2sin^2x. Also, cos^2x - sin^2x => cos^2x - (1-cos^2x) => 2cos^2x -1 = cos 2x. Hence cos^2x - sin^2x also equals cos 2x.

OpenStudy (anonymous):

So yes then...?

OpenStudy (anonymous):

lol the sum is done!? of course cos 2x = cos 2x. :P

OpenStudy (anonymous):

That's what I thought ha. Thanks